What is the velocity o π- meson whose proper means life time is 2.5× 10⅞ second

1
Expert's answer
2022-01-10T14:23:21-0500

We know that

"\tau=\frac{\tau_0}{\sqrt{(1-\frac{v^2}{c^2}})}"

"\tau=\tau_0\times2.5\times10^{\frac{7}{8}}"

"\tau_0\times2.5\times10^{\frac{7}{8}}=\frac{\tau_0}{\sqrt{(1-\frac{v^2}{c^2}})}"

Both sides take square

"6.25\times10^{\frac{7}{4}}=\frac{1}{{1-\frac{v^2}{c^2}}}"

"{{1-\frac{v^2}{c^2}}}=.16\times10^{\frac{-7}{4}}"

"{{1-\frac{v^2}{c^2}}}=2.845\times10^{-3}"

"\frac{v^2}{c^2}=0.9971"

"\frac{v}{c}=0.9985"

"v=0.9985c"